OpenStudy (anonymous):

Which expressions are equal to 8 when x = 6 and y = 2? Choose all answers that are correct. A. x – y + 4 B. y + x – 4 C. xy – 2 D. 5 + (x ÷ y)

OpenStudy (anonymous):

OK lets see

OpenStudy (anonymous):

copy and past the question again please and remove the extra space it makes it harder to work on this cause i gotta scroll up and down please

OpenStudy (anonymous):

A is right

OpenStudy (anonymous):

k

OpenStudy (anonymous):

nvm XD

OpenStudy (anonymous):

btw @mastermindkakashi your supposed to help them out not give answers i used to do that and got blocked so watch out for tha t

OpenStudy (anonymous):

so just try plugging in the numbers you know, for A. 6-2 is 4 and 4+4 is 8 so A. is correct

OpenStudy (anonymous):

6-2=4+4=8

OpenStudy (anonymous):

ITS A MUTIPLE CHOICE

OpenStudy (anonymous):

you talking to me>?

OpenStudy (anonymous):

2+6=8+4=12 X

OpenStudy (anonymous):

Let me see myself k hmm...

OpenStudy (anonymous):

6x2=12-2=10

OpenStudy (anonymous):

they already said the answer is the first one A

OpenStudy (anonymous):

then do that for the next one. 2 + 6 – 4. that is not correct because that equals 4

OpenStudy (anonymous):

6/2=3+5=8 A and D are correct

OpenStudy (anonymous):

there are two answers. the last one is also correct

OpenStudy (anonymous):

also we should be giving the answers directly, we can onlt give them is they find out themselves OR we work it out

OpenStudy (anonymous):

OK i think its A and D

OpenStudy (anonymous):

yup good job :D

OpenStudy (anonymous):

THANKS K NEXT QUESTION

OpenStudy (anonymous):

Evaluate the expression when x = 12 and y = 2. A. B. 2 C. D. 8

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

please post it as a new question next time. :)

OpenStudy (anonymous):

i will do that now k just a sec click